OpenStudy (anonymous):

Evaluate the integral for x > 1/2: integral x^2/sqrt(36x^2-9)

OpenStudy (dumbcow):

Let \[x = \frac{\sec u}{2}\] \[dx = \frac{\sec u \tan u}{2}\] Note sec^2 -1 = tan^2 \[\rightarrow \frac{1}{24} \int\limits \frac{ \sec^{2} u}{\tan u} (\sec u \tan u) du\]

OpenStudy (anonymous):

is that the answer or do I have to take the integral of that?

OpenStudy (dumbcow):

no sorry i didnt finish, i just got you started by doing the substitution you still need to integrate